OpenStudy (anonymous):

If two lines intersect at a 90 degree angle, what can you say about the lines?

OpenStudy (anonymous):

You can say that the lines are perpendicular :D

OpenStudy (anonymous):

thank you! DO you know thus one : The rules of parallel and perpendicular lines are different in spherical geometry than in Euclidean geometry. true or false?

OpenStudy (anonymous):

Its true

OpenStudy (anonymous):

The image below shows two parallel lines cut by a transversal. What is the value of a?

OpenStudy (anonymous):

What you are going to do is (2a +3)º + aº = 180º

OpenStudy (anonymous):

thank you, & how to do this one? If two parallel lines are cut by a transversal and two alternate exterior angles measure 5x + 3 and 2x + 21, what is the value of x?

OpenStudy (anonymous):

So for 2a + 3 + a = 180 x=30??

OpenStudy (anonymous):

It would be a = 59, and for the other one since there alternate exterior angles they are going to equal each other. So you would set it up like this: 5x + 3 = 2x+ 21

OpenStudy (anonymous):

I have to go now but the answer is for the second one is x = 6, I will be on later to answer any other questions =D

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

I am back if you have anymore questions :D

OpenStudy (anonymous):

The image below shows two parallel lines cut by a transversal. What is the value of x? x = 70 right?

OpenStudy (anonymous):

Yep, perfect!

OpenStudy (anonymous):

okay, & how to solve this? What is the value of x?

OpenStudy (anonymous):

its means they are equal right? so it would be 2?

OpenStudy (anonymous):

These are vertical angles, which means they equal each other. You would set up the problem like this: 12x - 7 = 4x + 9

OpenStudy (anonymous):

Yep, thats correct

OpenStudy (anonymous):

okay & If lines are perpendicular, what do you know about the angles at which they intersect?

OpenStudy (anonymous):

okay, so if we have two lines that are intersecting we know that all the angles will be = to 90º, or right angles

OpenStudy (anonymous):

alright,The point where two rays meet to form an angle is called what?

OpenStudy (anonymous):

The vertex

OpenStudy (anonymous):

What is the value of x?

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

what is the value of x?

OpenStudy (anonymous):

also whats the value of x for this one 2? How do you find the value. I dont get how too?

OpenStudy (anonymous):

It is going to equal 4, for the first one, so it'll be set up like this: -3x+1=4, and for the second problem it'll be 4x-2=6. We get this because there are 6 spaces in between the two lines, does that make sense?

OpenStudy (anonymous):

oh okay i get it thanks. Is a segment bisector makes two segments that have equal length. true or false?

OpenStudy (anonymous):

True

OpenStudy (anonymous):

Last question : Which of the following is not a correct name for this line?

OpenStudy (anonymous):

what are the following options?

OpenStudy (anonymous):

sorry my bad they are:

OpenStudy (anonymous):

This either BC or C, I am not sure which one

OpenStudy (anonymous):

alright thank you!(:

OpenStudy (anonymous):

No problem, if you ever have anymore question you can send them to me